# High Performance Distributed Control Law for Large-Scale Linear Systems: A Partitioned Distributed Observer Approach

### v3

@@ III-A Network partition
From this example, we can see that if we design partitioned distributed observers based on the partitioning results shown in Figure 4 and 5, we can significantly reduce the dimension of observers on each agent. However, we must emphasize an important fact: the degree of reduction in observer dimension on agents is closely related to the similarity between the communication network and the physical network. Due to the arbitrariness of communication and physical networks, some strange partitions may occur. For example, two nodes that are neighbors in a physical network may be far apart in a communication network, which can lead to the occurrence of huge partitions. In such cases, the reduction in the dimension of observers on each agent will not be as pronounced as in Figure 4 and 5. Therefore, we introduce the concept of “network similarity”—defined by $S_{pc}=2|\mathcal{E}_{c}\cap\mathcal{E}_{p}|/(|\mathcal{E}_{c}|+|\mathcal{E}_{p}|)$—to improve the rigor when describing simulation results. In other words, when describing how much our partition algorithm reduces the dimension of observers on each agent, we must indicate the degree of network similarity because the lower the network similarity, the higher the possibility of large abnormal partitions, and the lower the possibility of significant reduction in observer dimension.
III-B Partitioned distributed observer
This subsection designs distributed observer for each partition based on the partition method in the previous subsection. Suppose that $\mathcal{O}_{p}$ is a partition belonging to $\mathscr{P}_{i}$, and then the state observer on the $i$th agent to itself takes the form of:
$\dot{\hat{x}}_{ii}^{(p)}=$ $A_{ii}\hat{x}_{ii}^{(p)}+\sum_{j\in\mathcal{N}_{i}}A_{ij}\bar{x}_{ij}+\theta\Gamma_{\epsilon}H_{i}\left(y_{i}-C_{i}\hat{x}_{ii}^{(p)}\right)$
$+B_{i}\hat{u}_{ii}+\gamma\theta^{n}\tilde{P}_{i}^{-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(\hat{x}_{ji}^{(p)}-\hat{x}_{ii}^{(p)}\right),$ (8)
where $\hat{x}_{ii}^{(p)}$ is the estimation of $x_{i}$ generated by observers on agent $i$ in partition $\mathcal{O}_{p}$; $\bar{x}_{ij}=1/N_{j,\mathscr{P}_{i}}\cdot\sum_{\mathcal{O}_{p}\in\mathscr{P}_{i}}\hat{x}_{ij}^{(p)}$ with $N_{j,\mathscr{P}_{i}}$ being the number of occurrences of $j$ in $\mathcal{O}(\mathscr{P}_{i})$; $\hat{x}_{ij}^{(p)}$ stands for the state estimation of the $j$th subsystem generated by the $i$th agent. $\Gamma_{\epsilon}=diag\{\epsilon^{n-1},\ldots,\epsilon,1\}$ is the high-gain matrix with $\epsilon=1/\theta$ and $\theta$ being the high-gain parameter; $H_{i}\in\mathbb{R}^{n\times p}$ and $\gamma$ are, respectively, the observer gain and coupling gain; $\tilde{P}_{i}=\Gamma_{\epsilon}P_{i}\Gamma_{\epsilon}^{-1}$ is the so-called weighted matrix and $P_{i}$ is a symmetric positive definite matrix that will be designed later; $\hat{u}_{ii}^{(p)}$ is the control input relying on $\bar{x}_{ij}$.
@@ III-B Partitioned distributed observer
$\dot{x}_{li}^{(p)}=$ $A_{ii}\hat{x}_{li}^{(p)}+\sum_{j\in\mathcal{O}(\mathscr{P}_{l})\cap\mathcal{N}_{i}}A_{ij}\bar{x}_{lj}$
$+B_{i}\hat{u}_{li}+\gamma\theta^{n}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(\hat{x}_{ji}^{(p)}-\hat{x}_{li}^{(p)}\right),$ (9)
where $\hat{x}_{li}^{(p)}$ is the estimation of $x_{i}$ generated by observers on agent $l$ in partition $\mathcal{O}_{p}$, and $\hat{u}_{li}$ is the control input relying on $\bar{x}_{lj}$.
@@
$\dot{e}_{ii}^{(p)}=$ $\left(A_{ii}-\theta\Gamma_{\epsilon}H_{i}C_{i}\right)e_{ii}^{(p)}+\sum_{j\in\mathcal{N}_{i}}A_{ij}\bar{e}_{ij}$
$+B_{i}(\hat{u}_{ii}-\bar{u}_{i})-\gamma\theta^{n}\tilde{P}_{i}^{-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(e_{ji}^{(p)}-e_{ii}^{(p)}\right),$ (10)
@@
$+B_{i}(\hat{u}_{li}-\bar{u}_{i})-\gamma\theta^{n}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(e_{ji}^{(p)}-e_{li}^{(p)}\right),$ (11)
@@
This section has achieved the first goal of this paper through designing the partition algorithm and partition distributed observer: the dimension of observers on agent $i$ ($\hat{x}_{i\star}\triangleq col\{\hat{x}_{ij}^{(p)},~{}j\in\mathcal{O}_{p},~{}\mathcal{O}_{p}\in\mathscr{P}_{i}\}$) is much smaller than that of the interconnected system.
@@
There are two differences between the design of the partitioned distributed observer and the traditional distributed observer [13, 15]. First, since the states of the $j$th subsystem may be repeatedly estimated by agent $i$ in multiple partitions, we introduce the fusion estimation ($\bar{x}_{ij}$) in the coupling part of observers on each agent. Second, agent $i$ needs to use the physical network neighbor information of agent $l$ when estimating the states of subsystem $l$ (see in (9)). In traditional distributed observers, this information can be directly found in observers on the $i$th agent. However, since the dimension of observers on the $i$th agent in this paper is much smaller than that of the interconnected system, its states fail to cover the neighbor states of subsystem $l$. Therefore, we need to eliminate the coupling relationships ($\sum_{j\in\mathcal{O}(\mathscr{P}_{l})\cap\mathcal{N}_{i}}A_{ij}\bar{x}_{lj}$ in (9)) that cannot be obtained by agent $i$ when designing $\hat{x}_{il}$. In other words, to reduce the dimension of observers on each agent, we have to artificially introduce model mismatch to compensate for the lack of information caused by observer dimension reduction. Moreover, the control input $\hat{u}_{il}$ in observers on the $i$ agent also needs the state estimation $\hat{x}_{l\star}$. So, multiple estimates and model mismatch also occur in the control input ($\hat{u}_{ii}$ and $\hat{u}_{li}$) of the partitioned distributed observer (8) and (9). The content introduced here is the novelty of the partitioned distributed observer design in this section, and it is also one of the main difficulties to be solved later.
@@ IV Distributed control law for large-scale linear systems
Section III has reduced the dimension of observers on each agent by partitioning the network. This section will analyze the performance of the partitioned distributed observer and the closed-loop system after the dimension reduction of the local observer. Subsection IV-A shows the design of distributed control law. The stability of the error dynamics (10) and (11) and the closed-loop system will be proved in subsection IV-B and IV-C, respectively. Subsection IV-D will show the achievement of the second goal of this paper.
@@ IV-A Distributed control law
State feedback control law can be employed in the large-scale system (3) and (4). By observing the dynamics of the $i$th subsystem (1) and (2), the globally asymptotically stable control law gives rise to
@@
where $K_{ij}$ is designed so that $A+BK$ is a Hurwitz matrix with $K=[K_{ij}]_{i,j=1}^{N}$. However, (12) is the state feedback control law with precise system states. Hence, one should replace them with the state estimations generated by the observers on the $i$th agent. In the case of the traditional distributed-observer-based distributed control law [2, 9, 10, 11], the exact states of (12) can be directly replaced by state estimations of the $i$th local observer. However, since this paper studies large-scale systems and does not want every agent to reconstruct all the states of the whole system, the concept of partition is introduced and thus leads to a fusion estimation problem (See Remark 4. For example, agent $31$ in the example of Figure 4 and Figure 5 has estimated its states three times in three partitions). Hence, $\bar{x}_{ii}=1/|\mathscr{P}_{i}|\cdot\sum_{\mathcal{O}_{p}\in\mathscr{P}_{i}}\hat{x}_{ii}^{(p)}$, and $\bar{x}_{li}=1/N_{i,\mathscr{P}_{l}}\cdot\sum_{\mathcal{O}_{p}\in\mathscr{P}_{l}}\hat{x}_{li}^{(p)}$ (note that $1/N_{i,\mathscr{P}_{i}}=1/|\mathscr{P}_{i}|$) are expected to be employed in the control law of the $i$th subsystem, i.e.,
@@
Control law (12) is with precise and global states and is thus regarded as the centralized control law. (13) and (15) are employed in distributed observer (8) and (9). As for (14), it is the distributed-observer-based distributed control. The closed-loop system combined with (3) and (14) is the main research object of this paper.
@@
In addition, a discussion about $K$ should be given. Designing the controller gain $K$ in this section requires the use of global information. This approach is to ensure a meaningful comparison between the performance of the distributed control law proposed in this paper and that of the centralized control law (at least, the form of the control law should be the same). In many practical systems, it is sufficient to achieve the performance of centralized control by designing the control gain in a distributed manner. For example, in problems such as unmanned vehicle formations and coordinated control of microgrids (including the majority of practical systems that high-order systems can represent), each agent can design its own control gain. Then, based on the information provided by the partitioned distributed observer, it is possible to achieve distributed control laws with centralized control performance.
IV-B Performance of partitioned distributed observer
This subsection mainly focuses on the stability of error dynamics of partitioned distributed observers. Before that, we need to provide an important lemma. Denote $e_{\star i}^{(p)}=col\{e_{ji}^{(p)},~{}j\in\mathcal{O}_{p}\}$ and then we can state the follows.
@@ IV-B Performance of partitioned distributed observer
Now, the main theorem of this subsection can be given. Note that this is a preliminary conclusion on the stability of error dynamics of partitioned distributed observers. The complete conclusion will be presented in Theorem 3 in the next subsection.
@@
1) $H_{i}$ is chosen as $\theta^{n-1}\bar{H}_{i}$ such that $\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}$ is a Hurwitz matrix (see more detail of $\bar{H}_{i}$ in Remark 6) for all $i=1,\ldots,N$ where $\bar{A}_{ii}=\frac{1}{\theta^{n}}\Gamma_{\epsilon}^{-1}A_{ii}\Gamma_{\epsilon}$ and $\bar{C}_{i}=C_{i}\Gamma_{\epsilon}$;
@@
$\gamma>\frac{1}{\theta\underline{\lambda}_{\mathcal{O}_{p}}}\left(\lambda_{A}/\theta^{n-1}+2\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)\right),$ (21)
where $\underline{\lambda}_{\mathcal{O}_{p}}=\min_{i=1,\ldots,N}\{\lambda_{\mathcal{O}_{p}},~{}\mathcal{O}_{p}\in\mathscr{P}_{i}\}$ with $\lambda_{\mathcal{O}_{p}}=\underline{\sigma}(\mathcal{L}_{\mathcal{O}_{p}}+\mathcal{S}_{\mathcal{O}_{p}})$, $\mathcal{L}_{\mathcal{O}_{p}}$ being the Laplacian matrix of the subgraph among $\mathcal{O}_{p}$ as well as $\mathcal{S}_{\mathcal{O}_{p}}=diag\{1,0,\ldots,0\}\in\mathbb{R}^{|\mathcal{O}|_{p}\times|\mathcal{O}|_{p}}$; $\lambda_{A}=\max_{i}\{\bar{\sigma}(A_{ii}+A_{ii}^{T})\}$, $\lambda_{P}=\max_{i}\{\bar{\sigma}(P_{i})\}$, $\bar{\lambda}=\max_{i}\{\lambda_{i}\}$ with $\lambda_{i}=|\mathscr{P}_{i}|\max\{|\mathcal{O}_{p}|,~{}p\in\mathscr{P}_{i}\}$, and $\rho=\max_{i}\{\rho_{i}\}$ with $\rho_{i}=\max\{\rho_{li},~{}l\in\mathcal{O}(\mathscr{P}_{i})\}$.
@@
Set $\eta_{ii}^{(p)}=\Gamma_{\epsilon}^{-1}e_{ii}^{(p)}$ and $\eta_{ji}^{(p)}=e_{ji}^{(p)}$, then, based on equation (10), the dynamics of $\eta_{ii}^{(p)}$ gives rise to
$\dot{\eta}_{ii}^{(p)}=$ $\theta^{n}\left(\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}\right)\eta_{ii}^{(p)}+\Gamma_{\epsilon}^{-1}\sum_{j\in\mathcal{N}_{i}}A_{ij}\bar{\eta}_{ij}$
$+\Gamma_{\epsilon}^{-1}B_{i}(\hat{u}_{ii}-\bar{u}_{i})-\gamma\theta^{n}P_{i}^{-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(\eta_{ji}^{(p)}-\eta_{ii}^{(p)}\right),$
where $\bar{\eta}_{ij}=\Gamma_{\epsilon}^{-1}\bar{e}_{ij}$. Denote $\eta_{\star i}^{(p)}=col\{\eta_{ji}^{(p)},~{}j\in\mathcal{O}_{p}\}$ and $\eta_{i\star}=col\{\eta_{ij},~{}j\in\mathcal{O}_{p},~{}\mathcal{O}_{p}\in\mathscr{P}_{i}\}$, then
@@
$-\gamma\theta^{n}\left(P_{\star i}^{(p)}\right)^{-1}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)},$
@@
$A_{\star i}^{(p)}=diag\left\{\theta^{n}(\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}),~{}\underbrace{A_{ii},\ldots,A_{ii}}_{|\mathcal{O}_{p}|-1}\right\},$
$\Gamma_{\star i}^{(p)}=diag\left\{\Gamma_{\epsilon}^{-1},~{}\underbrace{I_{n},\ldots,I_{n}}_{|\mathcal{O}_{p}|-1}\right\},$
@@
$+\theta^{n-1}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left(2P_{\star i}^{(p)}\Delta_{\star i}^{(p)}\eta_{\star i}^{(p)}+2P_{\star i}^{(p)}U_{\star i}^{(p)}\eta_{\star i}^{(p)}\right)$
$-\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\gamma\theta^{n}\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$\leq$ $\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\theta^{n}\left(\eta_{ii}^{(p)}\right)^{T}sym\left\{P_{i}\left(\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}\right)\right\}\eta_{ii}^{(p)}$
$-\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\gamma\theta^{n}\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
@@
$+2\theta^{n-1}\bar{\sigma}(P_{i})\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left(\sum_{l\in\mathcal{O}_{p}}\|A_{i}\|\left(\left\|\eta_{l\star}\right\|+W_{T_{1}}\right)\right.$
@@
$\left(\eta_{ii}^{(p)}\right)^{T}sym\left\{P_{i}\Lambda_{i}\right\}\eta_{ii}^{(p)}-\gamma\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$=$ $-\gamma\left(\eta_{ii}^{(p)}\right)^{T}\eta_{ii}^{(p)}-\gamma\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$=$ $-\gamma\left(\eta_{\star i}^{(p)}\right)^{T}\left(\left(\mathcal{L}_{\mathcal{O}_{p}}+\mathcal{S}_{\mathcal{O}_{p}}\right)\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$=$ $-\gamma\lambda_{\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2},$ (25)

### v4

@@ III-A Cover sets selecting
From this example, we can see that if we design cover-based distributed observers based on the coverage shown in Figure 4 and 5, we can significantly reduce the dimension of observers on each agent. However, we must emphasize an important fact: the degree of reduction in observer dimension on agents is closely related to the similarity between the communication network and the physical network. Due to the arbitrariness of communication and physical networks, some strange coverage may occur. For example, two nodes that are neighbors in a physical network may be far apart in a communication network, which can lead to the occurrence of huge cover set. In such cases, the reduction in the dimension of observers on each agent will not be as pronounced as in Figure 4 and 5. Therefore, we introduce the concept of “network similarity”—defined by $S_{pc}=2|\mathcal{E}_{c}\cap\mathcal{E}_{p}|/(|\mathcal{E}_{c}|+|\mathcal{E}_{p}|)$—to improve the rigor when describing simulation results. In other words, when describing how much our algorithm reduces the dimension of observers on each agent, we must indicate the degree of network similarity because the lower the network similarity, the higher the possibility of large abnormal cover sets, and the lower the possibility of significant reduction in observer dimension.
III-B Cover-based distributed observer
This subsection designs distributed observer for each cover set based on the coverage solving method in the previous subsection. Suppose that $\mathcal{O}_{p}$ is a cover set belonging to $\mathscr{P}_{i}$, and then the state observer on the $i$th agent to itself takes the form of:
$\dot{\hat{x}}_{ii}^{(p)}=$ $A_{ii}\hat{x}_{ii}^{(p)}+\sum_{j\in\mathcal{N}_{i}}A_{ij}\bar{x}_{ij}+\theta\Gamma_{\theta}^{-1}H_{i}\left(y_{i}-C_{i}\hat{x}_{ii}^{(p)}\right)$
$+B_{i}\hat{u}_{ii}+\gamma\theta^{n-1}\tilde{P}_{i}^{-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(\hat{x}_{ji}^{(p)}-\hat{x}_{ii}^{(p)}\right),$ (8)
where $\hat{x}_{ii}^{(p)}$ is the estimation of $x_{i}$ generated by observers on agent $i$ in cover set $\mathcal{O}_{p}$; $\bar{x}_{ij}=1/N_{j,\mathscr{P}_{i}}\cdot\sum_{p,~\mathcal{O}_{p}\in\mathscr{P}_{i}}\hat{x}_{ij}^{(p)}$ with $N_{j,\mathscr{P}_{i}}$ being the number of occurrences of $j$ in $\mathcal{O}(\mathscr{P}_{i})$; $\hat{x}_{ij}^{(p)}$ stands for the state estimation of the $j$th subsystem generated by the $i$th agent. $\Gamma_{\theta}=diag\{\theta^{n-1},\ldots,\theta,1\}$ is the high-gain matrix with $\theta$ being the high-gain parameter; $H_{i}\in\mathbb{R}^{n\times p}$ and $\gamma$ are, respectively, the observer gain and coupling gain; $\tilde{P}_{i}=\Gamma_{\epsilon}P_{i}\Gamma_{\epsilon}^{-1}$ is the so-called weighted matrix and $P_{i}$ is a symmetric positive definite matrix that will be designed later; $\hat{u}_{ii}^{(p)}$ is the control input relying on $\bar{x}_{ij}$.
@@ III-B Cover-based distributed observer
$\dot{\hat{x}}_{li}^{(p)}=$ $A_{ii}\hat{x}_{li}^{(p)}+\sum_{j\in\mathcal{O}(\mathscr{P}_{l})\cap\mathcal{N}_{i}}A_{ij}\bar{x}_{lj}$
$+B_{i}\hat{u}_{li}+\gamma\theta^{n-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(\hat{x}_{ji}^{(p)}-\hat{x}_{li}^{(p)}\right),$ (9)
where $\hat{x}_{li}^{(p)}$ is the estimation of $x_{i}$ generated by observers on agent $l$ in cover set $\mathcal{O}_{p}$, and $\hat{u}_{li}$ is the control input relying on $\bar{x}_{lj}$.
@@
$\dot{e}_{ii}^{(p)}=$ $\left(A_{ii}-\theta\Gamma_{\theta}^{-1}H_{i}C_{i}\right)e_{ii}^{(p)}+\sum_{j\in\mathcal{N}_{i}}A_{ij}\bar{e}_{ij}$
$+$ $B_{i}(\hat{u}_{ii}-\bar{u}_{i})-\gamma\theta^{n-1}\tilde{P}_{i}^{-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(e_{ji}^{(p)}-e_{ii}^{(p)}\right),$ (10)
@@
$+$ $B_{i}(\hat{u}_{li}-\bar{u}_{i})-\gamma\theta^{n-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(e_{ji}^{(p)}-e_{li}^{(p)}\right),$ (11)
@@
This section has achieved the first goal of this paper through designing the coverage solving algorithm and cover-based distributed observer: the dimension of observers on agent $i$ ($\hat{x}_{i\star}\triangleq col\{\hat{x}_{ij}^{(p)},~j\in\mathcal{O}_{p},~\mathcal{O}_{p}\in\mathscr{P}_{i}\}$) is much smaller than that of the interconnected system.
@@
There are two differences between the design of the cover-based distributed observer and the traditional distributed observer [7, 9]. First, since the states of the $j$th subsystem may be repeatedly estimated by agent $i$ in multiple cover sets, we introduce the fusion estimation ($\bar{x}_{ij}$) in the coupling part of observers on each agent. Second, agent $i$ needs to use the physical network neighbor information of agent $l$ when estimating the states of subsystem $l$ (see in (9)). In traditional distributed observers, this information can be directly found in observers on the $i$th agent. However, since the dimension of observers on the $i$th agent in this paper is much smaller than that of the interconnected system, its states fail to cover the neighbor states of subsystem $l$. Therefore, we need to eliminate the coupling relationships ($\sum_{j\in\mathcal{O}(\mathscr{P}_{l})\cap\mathcal{N}_{i}}A_{ij}\bar{x}_{lj}$ in (9)) that cannot be obtained by agent $i$ when designing $\hat{x}_{il}$. In other words, to reduce the dimension of observers on each agent, we have to artificially introduce model mismatch to compensate for the lack of information caused by observer dimension reduction. Moreover, the control input $\hat{u}_{il}$ in observers on the $i$ agent also needs the state estimation $\hat{x}_{l\star}$. So, multiple estimates and model mismatch also occur in the control input ($\hat{u}_{ii}$ and $\hat{u}_{li}$) of the cover-based distributed observer (8) and (9). The content introduced here is the novelty of the cover-based distributed observer design in this section, and it is also one of the main difficulties to be solved later.
@@ IV Distributed control law for large-scale linear systems
Section III has reduced the dimension of observers on each agent by find a coverage of all agents. This section will analyze the performance of the cover-based distributed observer and the closed-loop system after the dimension reduction of the local observer. Subsection IV-A shows the design of distributed control law. The stability of the error dynamics (10) and (11) and the closed-loop system will be proved in subsection IV-B and IV-C, respectively. Subsection IV-D will show the achievement of the second goal of this paper.
@@ IV-A Distributed control law
A state feedback control law can be designed for the large-scale system described by (3) and (4). Considering the dynamics of the $i$-th subsystem in (1) and (2), a globally asymptotically stable control law takes the form:
@@
where the gain matrices $K_{ij}$ are designed using the LQR method, with $K=[K_{ij}]_{i,j=1}^{N}$ representing the complete centralized optimal gain matrix. For network systems where the dynamics and cost matrices exhibit a Spatially-Exponential Decaying (SED) structure [36], the optimal gain matrix $K$ possesses a quasi-SED property (or Quasi-Sparsity). Specifically, the norm of the coupling gain $\|K_{ij}\|$ decays exponentially with the distance between subsystem $i$ and $j$, $\text{dist}(i,j)$. Thus, the terms for $j\notin\mathcal{N}_{i}$ are indeed sufficiently small. More importantly, theoretical analyses, such as those in [36], demonstrate that when employing a truncated local controller as (12) (only use physical neighbor’s states), the performance gap between this distributed control law and the global optimal centralized control is exponentially small in $2$. Therefore, (12) can be regarded as an approximate centralized control law.
However, (12) is the state feedback control law with precise system states. Hence, one should replace them with the state estimations generated by the observers on the $i$th agent. In the case of the traditional distributed-observer-based distributed control law [2, 21, 22, 23], the exact states of (12) can be directly replaced by state estimations of the $i$th local observer. However, the concept of coverage is introduced and thus leads to a fusion estimation problem (See Remark 4. For example, agent $31$ in the example of Figure 4 and Figure 5 has estimated its states three times in three cover sets). Hence, $\bar{x}_{ii}=1/|\mathscr{P}_{i}|\cdot\sum_{\mathcal{O}_{p}\in\mathscr{P}_{i}}\hat{x}_{ii}^{(p)}$, and $\bar{x}_{li}=1/N_{i,\mathscr{P}_{l}}\cdot\sum_{\mathcal{O}_{p}\in\mathscr{P}_{l}}\hat{x}_{li}^{(p)}$ (note that $1/N_{i,\mathscr{P}_{i}}=1/|\mathscr{P}_{i}|$) are expected to be employed in the control law of the $i$th subsystem, i.e.,
@@
Control law (12) is with precise and global states and is thus regarded as the quasi centralized control law, which has the almost same performance of centralized control law. (13) and (15) are employed in distributed observer (8) and (9). As for (14), it is the distributed-observer-based distributed control. The closed-loop system combined with (3) and (14) is the main research object of this paper.
@@
IV-B Performance of cover-based distributed observer
This subsection mainly focuses on the stability of error dynamics of cover-based distributed observers. Before that, we need to provide an important lemma. Denote $e_{\star i}^{(p)}=col\{e_{ji}^{(p)},~j\in\mathcal{O}_{p}\}$ and then we can state the follows.
@@ IV-B Performance of cover-based distributed observer
Now, the main theorem of this subsection can be given. Note that this is a preliminary conclusion on the stability of error dynamics of cover-based distributed observers. The complete conclusion will be presented in Theorem 3 in the next subsection.
@@
1) $H_{i}$ is chosen as $\theta^{n-2}\bar{H}_{i}$ such that $\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}$ is a Hurwitz matrix (see more detail of $\bar{H}_{i}$ in Remark LABEL:H) for all $i=1,\ldots,N$ where $\bar{A}_{ii}=\frac{1}{\theta^{n-1}}\Gamma_{\theta}A_{ii}\Gamma_{\theta}^{-1}$ and $\bar{C}_{i}=C_{i}\Gamma_{\theta}^{-1}$;
@@
$\gamma>\frac{1}{2\theta^{n-1}\underline{\lambda}_{\mathcal{O}_{p}}}\left(\lambda_{A}+2\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)\right),$ (21)
where $\underline{\lambda}_{\mathcal{O}_{p}}=\min_{i=1,\ldots,N}\{\lambda_{\mathcal{O}_{p}},~\mathcal{O}_{p}\in\mathscr{P}_{i}\}$ with $\lambda_{\mathcal{O}_{p}}=\underline{\sigma}(\mathcal{L}_{\mathcal{O}_{p}}+\mathcal{S}_{\mathcal{O}_{p}})$, $\mathcal{L}_{\mathcal{O}_{p}}$ being the Laplacian matrix of the subgraph among $\mathcal{O}_{p}$ as well as $\mathcal{S}_{\mathcal{O}_{p}}=diag\{1,0,\ldots,0\}\in\mathbb{R}^{|\mathcal{O}|_{p}\times|\mathcal{O}|_{p}}$; $\lambda_{A_{i}}=\max_{i}\{\bar{\sigma}(A_{ii}+A_{ii}^{T})\}$, $\lambda_{P}=\max_{i}\{\bar{\sigma}(P_{i})\}$, $\bar{\lambda}=\max_{i}\{\lambda_{i}\}$ with $\lambda_{i}=|\mathscr{P}_{i}|\max\{|\mathcal{O}_{p}|,~p\in\mathscr{P}_{i}\}$, and $\rho=\max_{i}\{\rho_{i}\}$ with $\rho_{i}=\max\{\rho_{li},~l\in\mathcal{O}(\mathscr{P}_{i})\}$.
@@
Set $\eta_{ii}^{(p)}=\Gamma_{\theta}e_{ii}^{(p)}$ and $\eta_{ji}^{(p)}=e_{ji}^{(p)}$, then, based on equation (10), the dynamics of $\eta_{ii}^{(p)}$ gives rise to
$\dot{\eta}_{ii}^{(p)}=$ $\theta^{n-1}\left(\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}\right)\eta_{ii}^{(p)}+\Gamma_{\theta}\sum_{j\in\mathcal{N}_{i}}A_{ij}\bar{\eta}_{ij}$
$+\Gamma_{\epsilon}^{-1}B_{i}(\hat{u}_{ii}-\bar{u}_{i})-\gamma\theta^{n-1}P_{i}^{-1}\sum_{j\in\mathcal{O}_{p}}\alpha_{ij}\left(\eta_{ji}^{(p)}-\eta_{ii}^{(p)}\right),$
where $\bar{\eta}_{ij}=\Gamma_{\theta}\bar{e}_{ij}$. Denote $\eta_{\star i}^{(p)}=col\{\eta_{ji}^{(p)},~j\in\mathcal{O}_{p}\}$ and $\eta_{i\star}=col\{\eta_{ij},~j\in\mathcal{O}_{p},~\mathcal{O}_{p}\in\mathscr{P}_{i}\}$, then
@@
$-\gamma\theta^{n-1}\left(P_{\star i}^{(p)}\right)^{-1}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)},$
@@
$A_{\star i}^{(p)}=diag\left\{\theta^{n-1}(\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}),~\underbrace{A_{ii},\ldots,A_{ii}}_{|\mathcal{O}_{p}|-1}\right\},$
$\Gamma_{\star i}^{(p)}=diag\left\{\Gamma_{\theta},~\underbrace{I_{n},\ldots,I_{n}}_{|\mathcal{O}_{p}|-1}\right\},$
@@
$+\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left(2P_{\star i}^{(p)}\Delta_{\star i}^{(p)}\eta_{\star i}^{(p)}+2P_{\star i}^{(p)}U_{\star i}^{(p)}\eta_{\star i}^{(p)}\right)$
$-2\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\gamma\theta^{n-1}\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$\leq$ $\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\theta^{n-1}\left(\eta_{ii}^{(p)}\right)^{T}sym\left\{P_{i}\left(\bar{A}_{ii}-\bar{H}_{i}\bar{C}_{i}\right)\right\}\eta_{ii}^{(p)}$
$-2\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\gamma\theta^{n-1}\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
@@
$+2\bar{\sigma}(P_{i})\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left(\sum_{l\in\mathcal{O}_{p}}\|A_{i}\|\left(\left\|\eta_{l\star}\right\|+W_{T_{1}}\right)\right$.
@@
$\left(\eta_{ii}^{(p)}\right)^{T}sym\left\{P_{i}\Lambda_{i}\right\}\eta_{ii}^{(p)}-2\gamma\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$=$ $-2\gamma\left(\eta_{ii}^{(p)}\right)^{T}\eta_{ii}^{(p)}-2\gamma\left(\eta_{\star i}^{(p)}\right)^{T}\left(\mathcal{L}_{\mathcal{O}_{p}}\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$=$ $-2\gamma\left(\eta_{\star i}^{(p)}\right)^{T}\left(\left(\mathcal{L}_{\mathcal{O}_{p}}+\mathcal{S}_{\mathcal{O}_{p}}\right)\otimes I_{n}\right)\eta_{\star i}^{(p)}$
$=$ $-2\gamma\lambda_{\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2},$ (25)